Solve the equation: x+6/x=-5
Answer by RAY100(1637)

(x+6) / x = -5
cross multiply
x+6 = -5x
6x +6 =0
6(x+1) =0
x=-1
check
(-1) +6 = -5 (-1)
5=5,,,,,,,ok